OpenStudy (anonymous):

Evaluate: [1 -6][1 4] [2 0][9 8]

OpenStudy (anonymous):

This is simple. Try yourself..:)

OpenStudy (anonymous):

do i multiply?

OpenStudy (anonymous):

Yes matrix multiplication.

OpenStudy (anonymous):

Can we carry out such an operation? Refer your previous question.

OpenStudy (anonymous):

1x1..-6x4..?

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ 4\end{matrix}\right)\left[\begin{matrix}1 & 6\end{matrix}\right] \]

OpenStudy (anonymous):

is the matrix are like this?

OpenStudy (anonymous):

For example, you may put the following answer down for a 2 x 2 matrix. row 1 column 1 = 4 row 1 column 2 = 5 row 2 column 1 = 0 row 2 column 2 = -1

OpenStudy (anonymous):

if it's a 1x2 * 2x1 matrix you can multiply as you said. In question you've shown like 1x2* 1x2 matrix.

OpenStudy (anonymous):

ok..can you hold on so i can show you what i got?

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

[1 -24] [18 0]

OpenStudy (anonymous):



OpenStudy (anonymous):

ok..?

OpenStudy (anonymous):

was i wrong?

OpenStudy (anonymous):

There is a difference between [1 -6][1 4] and what I've shown in the fig. In the fig it's 1x2 * 2x1 matrix and what you've written in your question is 1x2*1x2 multiplication. Which is not possible. If the matrix are 1x2* 2x1 then answwer to the first question is [-24]. Hope it helps.

OpenStudy (anonymous):

For further info you can refer this web.. http://www.mathwarehouse.com/algebra/matrix/multiply-matrix.php

OpenStudy (anonymous):

you can note this is a linear combination 1 times firs row,-6 times second rown

hero (hero):

I was trying to post some matrices earlier but failed